OpenStudy (anonymous):

the length of a rectangle is 3 times the width. the perimeter is 96 cm. find the width and length.

OpenStudy (anonymous):

L = 3W 2L + 2W = 96 Solve simultaneously and you get l=36, w=12

OpenStudy (anonymous):

sheg how did you get that

OpenStudy (anonymous):

Sheg's answer clearly isn't right since the perimeter of such a figure is not 96

OpenStudy (anonymous):

moneybird's answer is also not right

OpenStudy (anonymous):

That is given

OpenStudy (anonymous):

yes they are wrong

OpenStudy (anonymous):

it should be 4w = 48 so w = 12 length = 36

OpenStudy (anonymous):

l=36, w=12 as I said in my original answer

OpenStudy (anonymous):

ktklown you are right how did u get that

OpenStudy (anonymous):

The length is 3 times the width, so L=3W The perimeter is the sum of the four sides, two lengths and two widths. 2L + 2W = 96 Then you solve simultaneously.

OpenStudy (anonymous):

Since L=3W, substitute that into the 2nd equation 2(3W) + 2W = 96 6W + 2W = 96 8W = 96 W = 12

OpenStudy (anonymous):

yeah in hurried ness i divided 96 by 16 instead of 8 so uit should be 96/8 = 12 so width = 12 length = 36

OpenStudy (anonymous):

how do u find the length

OpenStudy (anonymous):

L=3W

OpenStudy (anonymous):

In general once you've solved for one variable in a system of two equations, just plug that value into either of the other two equations (whichever looks easier)

OpenStudy (anonymous):

You have two equations, L=3W and 2L+2W=96. Once you have W you just plug the value for W into either one of those equations and solve.

OpenStudy (anonymous):

so can you please solve 2l+ 2w= 96 please

OpenStudy (anonymous):

What part are you having trouble with

OpenStudy (anonymous):

finding the lenght

OpenStudy (anonymous):

So substitute the width into either equation. The other equation is probably easier.

OpenStudy (anonymous):

can u please show me how it is supposed to look please

OpenStudy (anonymous):

Take the equation you had, L=3W and substitute in W=12

OpenStudy (anonymous):

can u just solve it for me cuz i dont understand

OpenStudy (anonymous):

do you know what "substituting" means?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

OK so can you substitute W=12 into the equation L=3W?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what do you get

OpenStudy (anonymous):

36

OpenStudy (anonymous):

thats the answer

OpenStudy (anonymous):

thank you